a particle starts from initial speed u and retardation ‘a’ to come to rest in time T. The time taken tio cover first half of the total path travelled is:
1)T/2^0.5
2)T(1-1/2^0.5)
3)T/2
4)3T/4

Grade:11

2 Answers

bharat bajaj IIT Delhi
9 years ago
Total path travelled from newtons third equation :v^2 - u^2 = 2asd = u^2/2aNow, to find out time to travel distance d/2.General equation :
s = ut + 1/2at^2
u^2/4a = ut - 1/2. a . t^2u^2/a = 4ut - 2a.t^22(at)^2 - 4uat + u^2 = 0Discriminant = 16(ua)^2 - 8(ua)^2 = 8(ua)^2t = 4ua - 2v2 ua/ 4a^2t = (2u- v2u)/2at = T(1 - 1/2^0.5)Option 2) is correctTHanks
Bharat BajajIIT Delhi
askiitians faculty
KAPIL MANDAL
9 years ago
Thank you sir for this answer. Thank you very much.